# Flattening Hierarchies with Policy Bootstrapping

### v1

@@ 6.2 Locomotion results
State-based locomotion: As a method designed for long-horizon reasoning, SAW excels in all variants of the state-based locomotion tasks. It scales particularly well to long horizons, exhibiting the best performance of $\mathbf{73}\%$ across all tasks in antmaze-giant-navigate and is the first method to achieve non-trivial success in humanoidmaze-giant-navigate, reaching $\mathbf{35}\%$ success compared to the previous state-of-the-art of $12\%$ [31]. We demonstrate that training subpolicies with subgoal representations scales poorly to the giant maze environments [Figure 3] but are critical to HIQL’s performance, emphasizing a fundamental tradeoff in hierarchical methods: subgoal representations are essential for making high-level policy prediction tractable, but those same representations can constrain policy expressiveness and limit overall performance. While other subgoal representation learning objectives may perform better than those derived from the value function, as in HIQL, this highlights the additional design complexity and tuning required for HRL methods. We also implement an offline variant of RIS [Appendix D] and find that it performs significantly worse than SAW with subgoal representations, which we suspect can be explained by our insights in Section 4.3.
@@ 7 Discussion
We presented Subgoal Advantage-Weighted Policy Bootstrapping (SAW), a simple yet effective policy extraction objective that leverages the subgoal structure of goal-conditioned tasks to scale to long-horizon tasks, without learning generative subgoal models. SAW consistently matches or surpasses current state-of-the-art methods across a wide variety of locomotion and manipulation tasks that require different timescales of control, whereas existing methods tend to specialize in particular task categories. Our method especially distinguishes itself in long-horizon reasoning, excelling in the most difficult locomotion tasks and scene-based manipulation. While the simplicity of our objective does introduce some practical limitations related to subgoal sampling, which we discuss in Appendix A, we find that avoiding explicit subgoal prediction is crucial for maintaining performance in large state spaces. By demonstrating a scalable approach to train unified policies for offline GCRL, we believe that SAW takes a step toward realizing the full potential of robotic foundation models in addressing the long-horizon, high-dimensional challenges of real-world control.
@@
Appendix A Limitations
A theoretical limitation of our approach, which is common to all hierarchical methods as well as RIS, occurs in our assumption that the optimal policy can be represented in the factored form $\pi^{\ell}(a\mid s,w)\pi^{h}(w\mid s,g)$. While this is true in theory (since we could trivially set $\pi^{h}(w\mid s,g)$ to a point distribution at $g$), practical algorithms typically fix the distance of the subgoals to a shorter distance of $k$ steps (or the midpoint in RIS), where subgoals $s_{t+k}$ are sampled from the future state distribution $p^{\mathcal{D}}_{\mathrm{traj}}(s_{t+k}\mid s_{t})$. Intuitively, if the dataset contains only suboptimal trajectories towards waypoints occurring $k$ steps later which are reachable in fewer than $k$ steps with an optimal low-level policy $\pi^{\ell}$, then the space of sampled subgoals will not contain these further state-subgoal pairs and any approximation of $\pi^{h}(w\mid s,g,U=1)$ (whether an explicit subgoal generator or our importance-weighted approach), will suffer additional approximation gaps.
However, as discussed in Section 4.3, we find that subgoals sampled from the future state distribution empirically work well with respect to goals also sampled from the future state distribution, which is common in practice [13, 12, 48, 10, 33]. However, we expect subgoal generator-based methods to have the edge when we sample from a goal distribution for which in-trajectory state-subgoal pairs tend to be highly suboptimal. While a generative subgoal model can synthesize “imagined” subgoals on which to bootstrap, our approach may require alternative subgoal-sampling strategies to reach the same level of performance.
Appendix B Planning Invariance
As an aside, we note that the discussions in this paper are closely related to the recently introduced concept of planning invariance [27], which describes a policy that takes similar actions when directed towards a goal as when directed towards an intermediate waypoint en route to that goal. In fact, we can say that subgoal-conditioned HRL methods achieve a form of planning invariance by construction, since they simply use the actions yielded by waypoint-conditioned policies to reach further goals. By minimizing the divergence between the full goal-conditioned policy and an associated subgoal-conditioned policy, both SAW and RIS can also be seen as implicitly enforcing planning invariance.
Appendix C Derivations of HIQL, RIS, and SAW Objectives
We cast the infinite-horizon, discounted GCRL formulation as an inference problem by constructing a probabilistic model via the likelihood function
@@ Appendix C Derivations of HIQL, RIS, and SAW Objectives
$p\left(U_{t}=1\mid\tau,\{w\},g\right)\propto\exp\left(\beta\sum_{t=0}^{\infty}\gamma^{t}A\left(s_{t},w_{t},g\right)\right),$
@@
where $\beta$ is an inverse temperature parameter and the binary variable $U$ can be intuitively understood as the event of reaching the goal $g$ as quickly as possible by passing through a subgoal $w$, or passing through a subgoal $w$ which is on the shortest path between $s_{t}$ and $g$.
We consider prior distributions $\pi$ of a factored hierarchical form
$p_{\pi}(\tau\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\,\pi^{\ell}(a\mid s_{t},w_{t})\,\pi^{h}(w_{t}\mid s_{t},g).$
C.1 HIQL derivation
Since HIQL learns two levels of a policy, we can use a variational posterior of the same form
$\pi_{\theta,\psi}(\tau\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\,\pi^{\ell}_{\theta}(a_{t}\mid s_{t},w_{t})\,\pi^{h}_{\psi}(w_{t}\mid s_{t},g).$
@@ C.1 HIQL derivation
$p\left(O_{t}=1\mid\tau,\{w\}\right)\propto\exp\left(\alpha\sum_{t=0}^{\infty}\gamma^{t}A\left(s_{t},a_{t},w_{t}\right)\right).$
With these definitions, we define the evidence lower bound (ELBO) on the joint optimality likelihood $p_{\pi}(O=1,U=1)$ for policy $\pi$
[…]
$=\log\mathbb{E}_{\pi_{\theta}(\tau\mid g),\pi_{\psi}(\{w\}\mid g),p(g)}\left[\frac{p(O=1,U=1,\tau,\{w\}\mid g)}{\pi_{\theta}(\tau\mid g)\pi_{\psi}(\{w\}\mid g)}\right]$
$\geq\mathbb{E}_{\pi_{\theta}(\tau\mid g),\pi_{\psi}(\{w\}\mid g),p(g)}\log\left[\frac{p(O=1,U=1,\tau,\{w\}\mid g)}{\pi_{\theta}(\tau\mid g)\pi_{\psi}(\{w\}\mid g)}\right]=\mathcal{J}(\theta,\psi).$
@@
We rewrite the discounted sum over time as an expectation over the (unnormalized) discounted stationary state distribution $\mu_{\pi}(s)=\sum_{t=0}^{\infty}\gamma^{t}p(s_{t}=s\mid\pi)$ induced by policy $\pi$. In practice, however, we optimize an approximation of $\mathcal{J}(\theta,\psi)$ by sampling from the dataset distribution over states $\mu_{\mathcal{D}}$. For brevity, we omit the conditionals in the expectations below, defining $\pi_{\theta}(a)\coloneq\pi_{\theta}(a\mid s,g)$ and $\pi_{\psi}(w)\coloneq\pi_{\psi}(w\mid s,g)$ in the expectations below
$\mathbb{E}_{\mu(s),\pi_{\theta}(a),\pi_{\psi}(w),p(g)}\left[\alpha A(s,a,w)+\log\left[\frac{\pi^{\ell}(a\mid s,w)}{\pi_{\theta}(a\mid s,w)}\right],\right]$
$\qquad+\mathbb{E}_{\mu(s),\pi_{\psi}(w),p(g)}\left[\beta A(s,w,g)+\log\left[\frac{\pi^{h}(w\mid s,g)}{\pi_{\psi}(w\mid s,g)}\right]\right]$
$=\mathbb{E}_{\mu(s),\pi_{\theta}(a),\pi_{\psi}(w),p(g)}\left[\alpha A(s,a,w)\right]-\mathbb{E}_{\mu(s),\pi_{\psi}(w),p(g)}\left[D_{\mathrm{KL}}\left[\pi_{\theta}(a\mid s,w)\|\pi^{\ell}(a\mid s,w)\right]\right]$
$\qquad+\mathbb{E}_{\mu(s),\pi_{\psi}(w),p(g)}\left[\beta A(s,w,g)\right]-\mathbb{E}_{\mu(s),p(g)}\left[D_{\mathrm{KL}}\left[\pi_{\psi}(w\mid s,g)\|\pi^{h}(w\mid s,g)\right]\right].$ (8)
HIQL separately optimizes the two summation terms, which correspond to the low- and high-level policies, respectively. Forming the Lagrangian with the normalization condition and solving for the optimal low- and high-level policies, as done in Abdolmaleki et al. [1] and Peng et al. [35], yields the HIQL AWR objectives:
$\mathcal{J}_{\pi^{\ell}}(\theta)$ $=\mathbb{E}_{\mu(s),\pi_{\theta}(a),\pi_{\psi}(w),p(g)}\left[\exp(\alpha A(s,a,w))\log\pi^{\ell}_{\theta}(a\mid s,w)\right]$
$\mathcal{J}_{\pi^{h}}(\psi)$ $=\mathbb{E}_{\mu(s),\pi_{\psi}(w),p(g)}\left[\exp(\beta A(s,w,g))\log\pi^{h}_{\psi}(w\mid s,g)\right].$
@@
While our derivation produces an on-policy expectation over actions and subgoals, the sampling distribution over states, actions, subgoals, and goals in the offline setting varies in practice (see Appendix C of Park et al. [31] for commonly used goal distributions).
C.2 RIS and SAW derivations

### v3

@@ 6.2 Locomotion results
State-based locomotion: As a method designed for long-horizon reasoning, SAW excels in all variants of the state-based locomotion tasks. It scales particularly well to long horizons, exhibiting the best performance of $\mathbf{73}\%$ across all tasks in antmaze-giant-navigate and is the first method to achieve non-trivial success in humanoidmaze-giant-navigate, reaching $\mathbf{35}\%$ success compared to the previous state-of-the-art of $12\%$ [park_ogbench_2024]. We demonstrate that training subpolicies with subgoal representations scales poorly to the giant maze environments [Figure 3] but are critical to HIQL’s performance, emphasizing a fundamental tradeoff in hierarchical methods: compact subgoal representations are essential for making high-level policy prediction tractable, but those same representations can constrain policy expressiveness and limit overall performance. While other subgoal representation learning objectives may perform better than those derived from the value function, as in HIQL, this highlights the additional design complexity and tuning required for HRL methods. We also implement an offline variant of RIS [Appendix C] and find that it performs significantly worse than SAW with subgoal representations, which we suspect can be explained by our insights in Section 4.3.
@@ 6.3 Manipulation results
7 Limitations
A theoretical limitation of our approach, which is common to all hierarchical methods as well as RIS, occurs in our assumption that the optimal policy can be represented in the factored form $\pi^{\ell}(a\mid s,w)\pi^{h}(w\mid s,g)$. While this is true in theory (since we could trivially set $\pi^{h}(w\mid s,g)$ to a point distribution at $g$), practical algorithms typically fix the distance of the subgoals to a shorter distance of $k$ steps, where subgoals $s_{t+k}$ are sampled from the distribution $p^{\mathcal{D}}_{\mathrm{traj}}(s_{t+k}\mid s_{t})$ induced by the behavior policy $\pi_{\beta}$. This introduces bias by limiting the space of subgoals to $k$-step future occupancy distribution of the behavior policy $\pi_{\beta}$, rather than that of the current learned policy. However, as discussed in Section 4.3, we find that subgoals sampled from the future state distribution work empirically well with respect to goals also sampled from the future state distribution, which is common in practice [gupta_relay_2020, ghosh_learning_2020, yang_rethinking_2021, eysenbach_contrastive_2022, park_hiql_2023].
We suspect that methods which generate subgoals may perform better in datasets that require a high degree of stitching: they can synthesize “imagined” subgoals on which to bootstrap, while our approach may require alternative sampling strategies to reach the same level of performance. While SAW maintains strong performance compared to baselines when trained on suboptimal stitch datasets, its performance does degrade in the highly suboptimal explore datasets [Appendix J].
@@ 8 Discussion
We presented Subgoal Advantage-Weighted Policy Bootstrapping (SAW), a simple yet effective policy extraction objective that leverages the subgoal structure of goal-conditioned tasks to scale to long-horizon tasks, without learning generative subgoal models. SAW consistently matches or surpasses current state-of-the-art methods across a wide variety of locomotion and manipulation tasks that require different timescales of control, whereas existing methods tend to specialize in particular task categories. Our method especially distinguishes itself in long-horizon reasoning, excelling in the most difficult locomotion tasks and scene-based manipulation.
In addition to providing a denser learning signal, one of the primary functions of value bootstrapping is to reduce the variance of the target estimates. Does this also apply to policy bootstrapping? We hypothesize that the answer is yes: when the value functions are noisy, as in long-horizon tasks, the advantage weight on a given action, and therefore the gradient of the policy objective, can be extremely high-variance over the course of training. On the other hand, our bootstrapping objective benefits from a stable action target provided by a subpolicy, which in turn enjoys clearer advantage signals [Section 4.3]. We believe that utilizing such target policies is a promising direction for improving the stability of long-horizon policy learning, especially when value functions are noisy.
As a brief addendum, we suggest that the family of policy bootstrapping algorithms can be seen as a practical instantiation of the chunking theory of learning from neuroscience, which suggests that mastering complex, sequential skills initially involves breaking them down into a series of chunks [ramkumar_chunking_2016, brown_repetition_2023]. Individual chunks are blended into smoother and more efficient movements with repeated practice, providing a mechanism for acquiring progressively longer and more complex skills that can again be used as primitive chunks for planning, in a virtuous cycle of improvement over the horizon.
@@
Appendix A Reinforcement Learning as Probabilistic Inference
In an effort to make this work more self-contained, we provide a short overview of the probabilistic inference perspective on reinforcement learning, which we use to derive the HIQL, RIS, and SAW objectives in Appendix B. The general structure of these derivations will be quite similar to those discussed later on and follows closely from abdolmaleki_maximum_2018, although we do deviate in some places for clarity and correctness. For a more thorough treatment, interested readers may refer to the tutorial by levine_reinforcement_2018. We break this introduction into two sections: (1) a general introduction to the RL as inference framework, (2) and a discussion of the modifications and assumptions made to apply these ideas to offline RL with a fixed stationary distribution $\mu^{\mathcal{D}}(s)$.
A.1 RL as inference
To view optimal control as a probabilistic inference problem, we introduce an optimality variable $O_{t}$, which takes the form of a binary random variable where $O_{t}=1$ if the decision at time step $t$ was optimal, and $O_{t}=0$ if not. We define the joint distribution of optimality variables in a trajectory $\tau$ as
$p(O=1\mid\tau)\propto\exp\left(\sum_{t=0}^{\infty}\gamma^{t}r_{t}/\alpha\right),$
where $\alpha$ is a temperature hyperparameter and $r_{t}=r(s_{t},a_{t})$. This has a natural interpretation: if we assume deterministic dynamics, the probability of a trajectory induced by the optimal policy is proportional to its exponentiated return. We can view this as a “soft” maximization over actions compared the hard $\operatorname*{arg\,max}$ in greedy Q-learning [schulman_equivalence_2018]. A natural objective is then to optimize the marginal likelihood of optimality for policy $\pi$. Unfortunately, this is a function of an expectation over the space of trajectories, which is intractable to compute. Instead, we can form the evidence lower bound (ELBO) by using Jensen’s inequality and introducing a variational approximation $q(\tau)$
$\log p_{\pi}(O=1)$ $=\log\int p_{\pi}(\tau)p(O=1\mid\tau)d\tau$
$\geq\int q(\tau)\left[\log p(O=1\mid\tau)+\log\frac{p_{\pi}(\tau)}{q(\tau)}\right]d\tau$
$=\mathbb{E}_{q}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}/\alpha\right]-D_{\mathrm{KL}}\left(q(\tau)\|p_{\pi}(\tau)\right)$
$=\mathcal{J}(q).$
Under the assumption that $p_{\pi}(\tau)=p\left(s_{0}\right)\prod_{t=0}^{\infty}p\left(s_{t+1}\mid s_{t},a_{t}\right)\pi\left(a_{t}\mid s_{t}\right)$, i.e., $\pi$ is Markovian, we consider a family of variational distributions that factor in the same way as $p_{\pi}(\tau)$
$q(\tau)=p\left(s_{0}\right)\prod_{t=0}^{\infty}p\left(s_{t+1}\mid s_{t},a_{t}\right)q\left(a_{t}\mid s_{t}\right).$
This choice of distribution allows us to rewrite the divergence between trajectories as a factored distribution over individual timesteps, where all terms not depending on $q$ cancel out and we move the weighting parameter $\alpha$ to the second term (which is equivalent for optimization purposes)
@@ A.1 RL as inference
$\mathcal{J}(q)$ $=\mathbb{E}_{q(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]-\mathbb{E}_{q(\tau)}\left[\alpha\log\left(\frac{q(\tau)}{p_{\pi}(\tau)}\right)\right]$
$=\mathbb{E}_{q(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]-\mathbb{E}_{q(\tau)}\left[\sum_{t=0}^{\infty}\alpha\log\left(\frac{q(a_{t}\mid s_{t})}{\pi(a_{t}\mid s_{t})}\right)\right]$
$=\mathbb{E}_{q(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}-\alpha\log\left(\frac{q(a_{t}\mid s_{t})}{\pi(a_{t}\mid s_{t})}\right)\right].$ (10)
Prior work [abdolmaleki_maximum_2018] additionally discounts the per-timestep divergence term, so that optimizing $\mathcal{J}$ with respect to $q$ is equivalent to solving the standard RL problem with an augmented reward $\tilde{r}_{t}=r_{t}-\alpha\log\frac{q(a_{t}\mid s_{t})}{\pi(a_{t}\mid s_{t})}$. We follow this convention as well in the following section.
A.2 Offline policy optimization
To optimize the above objective, we define the state-action value associated with Equation 10 as
$Q^{\pi}(s,a)=r_{0}+\mathbb{E}_{q(\tau),s_{0}=s,a_{0}=a}\left[\sum_{t=1}^{\infty}\gamma^{t}\left[r_{t}-\alpha\log\frac{q(a_{t}\mid s_{t})}{\pi(a_{t}\mid s_{t})}\right]\right].$ (11)
We then perform a partial maximization of $\mathcal{J}$. The key difference between this and standard policy improvement, which updates $\pi(s)\leftarrow\operatorname*{arg\,max}_{a}r(s,a)+\gamma\mathbb{E}_{p(s^{\prime}\mid s,a)}\left[V\left(s^{\prime}\right)\right]$, is that we select the soft-optimal action instead of a hard $\operatorname*{arg\,max}$ and use the value function $V^{\pi}$ associated with the fixed policy $\pi$ instead of the current policy $q$.
We first expand $Q^{\pi}$ using the regularized Bellman operator, which replaces $r$ with $\tilde{r}$
$T^{\pi,q}$ $=\mathbb{E}_{q(a\mid s)}\left[r(s,a)-\alpha\log\frac{q(a\mid s)}{\pi(a\mid s)}+\gamma\mathbb{E}_{p\left(s^{\prime}\mid s,a\right)}\left[V^{\pi}\left(s^{\prime}\right)\right]\right]$
$=\mathbb{E}_{q(a\mid s)}\left[r(s,a)+\gamma\mathbb{E}_{p\left(s^{\prime}\mid s,a\right)}\left[V^{\pi}\left(s^{\prime}\right)\right]\right]-\alpha D_{\mathrm{KL}}\left(q(a\mid s)\|\pi(a\mid s)\right)$
$=\mathbb{E}_{q(a\mid s)}\left[Q^{\pi}(s,a)\right]-\alpha D_{\mathrm{KL}}\left(q(a\mid s)\|\pi(a\mid s)\right).$
This gives us a “one-step” KL-regularized objective
$\max_{q}\bar{\mathcal{J}}(q)$ $=\max_{q}\mathbb{E}_{\mu^{\pi}(s)}\left[T^{\pi,q}Q^{\pi}(s,a)\right]$
$=\max_{q}\mathbb{E}_{\mu^{\pi}(s)}\left[\mathbb{E}_{q(a\mid s)}\left[Q^{\pi}(s,a)\right]-\alpha D_{\mathrm{KL}}\left(q(a\mid s)\|\pi(a\mid s)\right)\right],$ (12)
where $\mu^{\pi}(s)$ is the discounted stationary state distribution induced by $\pi$.
This does not fully optimize $\mathcal{J}$ since we ignore the dependence of the Q-value on policy $q$. However, we can also learn $Q$ from off-policy data: in this work, we use a variant of implicit Q-learning [Equation 2], which learns a value function for an improved (implicit) policy via value iteration. While the soft-optimal actor learned via this optimization procedure may differ from the implicit actor of IQL, such decoupled approaches work empirically well in practice.
For mathematical convenience, we can turn the KL regularization term into a hard constraint
$\max_{q}\mathbb{E}_{\mu(s)}\left[\mathbb{E}_{q(a\mid s)}\left[Q^{\pi}(s,a)\right]\right]$
$\text{ s.t. }\mathbb{E}_{\mu(s)}\left[D_{\mathrm{KL}}\left(q(a\mid s),\pi\left(a\mid s\right)\right)\right]<\epsilon,$ (13)
and use Lagrangian duality to find a closed form for the optimal posterior
$q(a\mid s)\propto\exp(Q(s,a)/\alpha)\pi_{\theta}(a\mid s).$ (14)
This can be interpreted as a Bayesian update, where the prior $\pi_{\theta}$ is updated by the likelihood of optimality $\exp(Q(s,a)/\alpha)$. Given a tractable way to estimate the optimality-conditioned posterior distribution from samples, we can now turn policy optimization into a variational inference problem.
Expectation maximization for policy learning: Framing RL as an inference problem allows us to use tools from approximate variational inference to estimate the optimal state-conditional distribution over actions. The expectation-maximization (EM) algorithm is one such tool that gives rise to a family of policy improvement algorithms [dayan_using_1996, abdolmaleki_maximum_2018, peng_advantage-weighted_2019], and can be described as “solving a sequence of probability matching problems, where $\theta$ is chosen at each step to match as best it can a fictitious distribution that is determined by the average rewards experienced on the previous step[dayan_using_1996].
More concretely, EM algorithms for RL are based an iterative two-step process, consisting of:
  1. 1.
    Expectation (E) step: We construct a nonparametric, sample-based posterior $q(a\mid s)\propto\exp(M(s,a)/\alpha)\pi_{\theta}(a\mid s)$, where $M(s,a)$ is some performance measure, e.g., the Q-function [abdolmaleki_maximum_2018] or the advantage function [peng_advantage-weighted_2019]. This is exactly the “fictitious” distribution referenced above. As mentioned above, we replace $\pi_{\theta}$ with the fixed behavior policy $\pi_{\beta}$ in the offline setting.
  2. 2.
    Maximization (M) step: In this step, we update our parameterized policy towards this fictitious target by minimizing the KL divergence between the two. In practice, we are rarely able to find the setting of $\theta$ that globally minimizes the divergence—and neither do we want to, since $q$ is only a noisy approximation of the true optimal posterior and can suffer from high sampling error. Instead, we simply take a few gradient steps on $\theta$ towards $q$ before resampling.
When $q$ is non-parametric as described above, we can directly substitute in our expression for $q$ [Equation 14] with a normalizing constant into the KL divergence and simplify, turning both steps into the single objective of Maximum a Posteriori Policy Optimization [abdolmaleki_maximum_2018, MPO]
$\mathcal{J}_{\mathrm{MPO}}(\theta)=\mathbb{E}_{s\sim\mu^{\mathcal{D}}(s),a\sim\pi_{\beta}(a\mid s)}\left[\exp(Q(s,a)/\alpha)\log\pi_{\theta}(a\mid s)\right].$ (15)
A similar procedure can be used on an objective that maximizes the expected improvement $\eta(\pi)=J(\pi_{\theta})-J(\pi_{\beta})$ over the behavior policy $\pi_{\beta}$, giving rise to an advantage-weighted objective [peng_advantage-weighted_2019]
$\mathcal{J}_{\mathrm{AWR}}(\theta)=\mathbb{E}_{s\sim\mu^{\mathcal{D}}(s),a\sim\pi_{\beta}(a\mid s)}\left[\exp(A(s,a)/\alpha)\log\pi_{\theta}(a\mid s)\right].$ (16)
As a point of clarification, in the online setting our replay buffer $\mathcal{D}$ is populated by fresh experience from $\pi_{\theta}$, which can be thought of as the prior distribution corresponding to the posterior $q$. However, in the offline setting, the prior is fixed to the offline dataset $\mathcal{D}$, and $\pi_{\theta}$ instead takes the role of a parametric estimate of the shape of $q$, that is repeatedly updated over sampled batches.
Appendix B Derivation and Unification of the HIQL, RIS, and SAW Objectives
As in the standard inference framework introduced in Appendix A, we cast the infinite-horizon, discounted GCRL formulation as an inference problem by constructing a probabilistic model. However, we introduce a new variable corresponding to subgoal optimality, via the likelihood function
$p\left(U=1\mid\tau,\{w\},g\right)\propto\exp\left(\beta\sum_{t=0}^{\infty}\gamma^{t}A\left(s_{t},w_{t},g\right)\right),$
@@ Appendix B Derivation and Unification of the HIQL, RIS, and SAW Objectives
where $\beta$ is an inverse temperature parameter and the binary variable $U$ can be intuitively understood as the event of reaching the goal $g$ as quickly as possible by passing through a subgoal $w$, or passing through a subgoal $w$ which is on the shortest path between $s_{t}$ and $g$. Following park_hiql_2023, we elect to use an advantage function to reduce variance, which is especially important when considering multi-step advantage estimates as we do here.
We consider policies $\pi^{\ell}$ and $\pi^{h}$ of a factored hierarchical form
$p_{\pi}(\tau\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\,\pi^{\ell}_{\theta}(a\mid s_{t},w_{t})\,\pi^{h}_{\psi}(w_{t}\mid s_{t},g),$
where the low-level policy over actions $\pi^{\ell}$ and high-level policy over subgoals $\pi^{h}$ are parameterized by $\theta$ and $\psi$, respectively. However, for the sake of clarity, we drop the parameters from the notation in the following derivations.
B.1 HIQL derivation
For HIQL’s hierarchical policy, we consider distributions $q^{\ell}$ and $q^{h}$ of the same form
$q(\tau\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\,q^{\ell}(a_{t}\mid s_{t},w_{t})\,q^{h}(w_{t}\mid s_{t},g).$
@@ B.1 HIQL derivation
$p\left(O=1\mid\tau,\{w\}\right)\propto\exp\left(\alpha\sum_{t=0}^{\infty}\gamma^{t}A\left(s_{t},a_{t},w_{t}\right)\right).$
With these definitions, we define the evidence lower bound (ELBO) on the joint optimality likelihood $p_{\pi}(O=1,U=1)$ for policy $\pi$ and posterior $q$ as
$\log p_{\pi}(O=1,U=1)$ $=\log\int p(g)\,p(O=1,U=1,\tau,\{w\}\mid g)d\{w\}\,d\tau\,dg$
$=\log\int p(g)q^{\ell}(\tau\mid g)q^{h}(\{w\}\mid g)\frac{p(O=1,U=1,\tau,\{w\}\mid g)}{q^{\ell}(\tau\mid g)\,q^{h}(\{w\}\mid g)}d\{w\}\,d\tau\,dg$
$=\log\mathbb{E}_{q^{\ell}(\tau\mid g),q^{h}(\{w\}\mid g),p(g)}\left[\frac{p(O=1,U=1,\tau,\{w\}\mid g)}{q^{\ell}(\tau\mid g)q^{h}(\{w\}\mid g)}\right]$
$\geq\mathbb{E}_{q^{\ell}(\tau\mid g),q^{h}(\{w\}\mid g),p(g)}\log\left[\frac{p(O=1,U=1,\tau,\{w\}\mid g)}{q^{\ell}(\tau\mid g)q^{h}(\{w\}\mid g)}\right]=\mathcal{J}(q,\pi).$
@@
We rewrite the discounted sum over time as an expectation over the (unnormalized) discounted stationary state distribution $\mu_{\pi}(s)=\sum_{t=0}^{\infty}\gamma^{t}p(s_{t}=s\mid\pi)$ induced by policy $\pi$. In practice, however, we optimize an approximation of $\mathcal{J}(q,\theta,\psi)$ by sampling from the dataset distribution over states $\mu_{\mathcal{D}}$, as justified in Appendix A. For brevity, we omit the conditionals in the expectations below, defining $q^{\ell}(a)\coloneq q^{\ell}(a\mid s,g)$ and $q^{h}(w)\coloneq q^{h}(w\mid s,g)$
$\mathbb{E}_{\mu(s),q^{\ell}(a),q^{h}(w),p(g)}\left[\alpha A(s,a,w)+\log\left[\frac{\pi^{\ell}(a\mid s,w)}{q^{\ell}(a\mid s,w)}\right]\right]$
$\qquad+\mathbb{E}_{\mu(s),q^{h}(w),p(g)}\left[\beta A(s,w,g)+\log\left[\frac{\pi^{h}(w\mid s,g)}{q^{h}(w\mid s,g)}\right]\right]$
$=\mathbb{E}_{\mu(s),q^{\ell}(a),q^{h}(w),p(g)}\left[\alpha A(s,a,w)\right]-\mathbb{E}_{\mu(s),q^{h}(w),p(g)}\left[D_{\mathrm{KL}}\left[q^{\ell}(a\mid s,w)\|\pi^{\ell}(a\mid s,w)\right]\right]$
$\qquad+\mathbb{E}_{\mu(s),q^{h}(w),p(g)}\left[\beta A(s,w,g)\right]-\mathbb{E}_{\mu(s),p(g)}\left[D_{\mathrm{KL}}\left[q^{h}(w\mid s,g)\|\pi^{h}(w\mid s,g)\right]\right].$ (17)
HIQL separately optimizes the two summation terms, which correspond to the low- and high-level policies, respectively. Following abdolmaleki_maximum_2018 and peng_advantage-weighted_2019, we turn each KL divergence penalty into a hard constraint $D_{\mathrm{KL}}(\cdot\|\cdot)\leq\epsilon$, form the Lagrangian, and solve for the optimal low- and high-level policies to yield the sample-based $q^{\ell}$ and $q^{h}$ distributions
$q^{\ell}(a\mid s,g)$ $\propto\pi^{\ell}(a\mid s,w)\exp(A(s,a,w))$
$q^{h}(a\mid s,g)$ $\propto\pi^{h}(w\mid s,g)\exp(A(s,w,g)).$
Minimizing the KL divergence between these sample-based estimates of the optimal posteriors and their respective parametric policies $\pi^{\ell}_{\theta}$ and $\pi^{h}_{\psi}$ yields the bilevel AWR policy extraction objectives for HIQL
$\mathcal{J}^{\ell}_{\mathrm{HIQL}}(\theta)$ $=\mathbb{E}_{\mu(s),q^{\ell}(a),q^{h}(w),p(g)}\left[\exp(A(s,a,w))\log\pi^{\ell}_{\theta}(a\mid s,w)\right]$
$\mathcal{J}^{h}_{\mathrm{HIQL}}(\psi)$ $=\mathbb{E}_{\mu(s),q^{h}(w),p(g)}\left[\exp(A(s,w,g))\log\pi^{h}_{\psi}(w\mid s,g)\right].$
@@
While our derivation produces an on-policy expectation over states, actions, and subgoals, we approximate this objective with samples drawn from the dataset distribution following the justification in Appendix A. The goal distribution is not known a priori (see Appendix C of park_ogbench_2024 for commonly used goal distributions).
B.2 RIS and SAW derivations